OpenStudy (anonymous):

Complete the square to form a perfect square trinomial. x^2+6x+

OpenStudy (anonymous):

two steps a) what is half of 6?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

b) what is \(3^2\)?

OpenStudy (anonymous):

6

OpenStudy (anonymous):

lol no

OpenStudy (anonymous):

\[2\times 3=6\\3^2\neq 6\]

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

Why is there a line threw the equal signs

jimthompson5910 (jim_thompson5910):

\(\Large 3^2\) is the same as 3 times 3

jimthompson5910 (jim_thompson5910):

the line through the equal sign means "not equal"

OpenStudy (anonymous):

So how do you complete the square to form a perfect square trinomial

jimthompson5910 (jim_thompson5910):

Fill in the blank \[\Large 3^2 = 3*3 = \underline{ \ \ \ \ }\]

OpenStudy (anonymous):

The eqations is x^2+6x+

jimthompson5910 (jim_thompson5910):

yes you already posted that at the top

OpenStudy (anonymous):

So after the plus sign I write 3^2=3*3=

jimthompson5910 (jim_thompson5910):

what is 3*3 equal to?

OpenStudy (anonymous):

3*3 isn't it 9

jimthompson5910 (jim_thompson5910):

yes

jimthompson5910 (jim_thompson5910):

so 9 goes in the blank

jimthompson5910 (jim_thompson5910):

x^2+6x+9 is a perfect square which factors to (x+3)^2

OpenStudy (anonymous):

Thank you

jimthompson5910 (jim_thompson5910):

np